OpenStudy (anonymous):

Does anyone know how to solve this: IF 2 + 5 + 5 = 51 3 + 7 + 3 = 61 5 + 4 + 3 = 91 2 + 7 + 3 = 31 THEN 6 + 7 + 3 = ?

OpenStudy (anonymous):

Wicked question, huh?

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

No clue...

OpenStudy (anonymous):

121

OpenStudy (anonymous):

How did you get 121?

OpenStudy (anonymous):

by using elimination 3+7+3=61 1+(2+7+3)=61 then 1= 61-31=30 then 6+7+3=121( 6 is 2*3 is 1*2=30*2=60